OpenStudy (mathlegend):

Logs.. solve the equation by expressing each side as a power of the same base and then equating exponents. (I have an answer but it is not an option)

OpenStudy (mathlegend):

\[64^{x}=\frac{ 1 }{ \sqrt{4} }\]

OpenStudy (mathlegend):

When I solved I got ... x = -1/12 Is this right or wrong?

OpenStudy (mathlegend):

Oh I figured it out... I used the wrong base... haha!